Juila12001:

does number 9 equal 18

Shadow:

I assume there is an image associated with this problem?

Juila12001:

yea

Shadow:

Can you take a screenshot or upload it?

Juila12001:

yea

Juila12001:



Juila12001:

https://ibb.co/mFAcmn idk if u can see the url

justasoftie:

So the measurements of angles add up to 360 degress I believe.

justasoftie:

Now what you would have to do is add it all up and find the missing angle measurement

justasoftie:

No I was wrong. They are up to 180 degrees. I'm really tired, I apologize

justasoftie:

Add*

justasoftie:

Wait I read your question wrong. You wanted your answer to be checked, correct?

Juila12001:

yea

justasoftie:

so you said it would be 18?

justasoftie:

This is really confusing me now

Juila12001:

all i did was cos; 46/48 and divide them and then on the chart it said 18 but I thinks thats wrong

Shadow:

It says the answer is 18?

Juila12001:

yea

Shadow:

Hmm, I get 16.6

Juila12001:

how

Shadow:

\[\cos \theta = \frac{ 46 }{ 48 }\] \[\theta = \cos^{-1} (\frac{ 46 }{ 48 })\] \[\theta = 16.6\]

Juila12001:

https://ibb.co/d0V2mn this the chart thing i used

Shadow:

Oh, I thought this was a study guide of some sort and 18 was the correct answer.

Juila12001:

ok

Shadow:

In the answer key

Shadow:

Why are you referring to that chart?

Juila12001:

our teacher gave it us

Shadow:

Are you allowed to use a calculator?

Juila12001:

yea

Juila12001:

can u help with 7

Shadow:

Mhm, that's how you would solve for #9 then. \[Cosine = \frac{ Adjacent }{ Hypotenuse }\] That's why I did \[\cos \theta = \frac{ 46 }{ 48 }\] 46 = adjacent side to the angle 48 = the hypotenuse

Shadow:

Sure, post a new question

Juila12001:

it was on the same picture question 7

Shadow:

1 attachment
Shadow:

This looks like Special Right Triangles. Does that sound familiar to you?

Juila12001:

yea

Shadow:

1 attachment
Juila12001:

do u use the 30 one?

Shadow:

Eh, actually, lets just do Pythag. We don't know for sure which one it is without an angle being given

Shadow:

\[(3)^2 + x^2 = (2 \sqrt 3)^2\] \[9 + x^2 = 4 \times 3 \] \[9 + x^2 = 12\] \[x^2 = 3\] \[x = \sqrt 3\]

Shadow:

It was a 30-60-90

Shadow:

lol

Shadow:

This is pretty easy though.

Juila12001:

how did get the answer 3

Shadow:

Subtracted 9 from both sides

Juila12001:

oh so the answer just 3

Shadow:

the answer is sqrt 3

Juila12001:

ok thx